A coil of wire 7 m long and with 220 turns carries a current of 1.6 A.            

  1. What is the magnitude of the magnetic field strength?                   


  1. Calculate the flux density if the coil is in a vacuum. 
  2. (Permeability of a vacuum is 1.257 × 10–6H/m.)                       


  1. Calculate the flux density inside a bar of carbon steel that is positioned within the coil. 




                                                     



1
Expert's answer
2021-09-29T18:35:06-0400

"Part(1)\\B=\frac{\mu_0 NI}{2\pi r}"

"B=\frac{4\pi \times220\times1.6}{2\pi \times7}"

"B=1.0057\times10^{-5}T"

Part(2)


"B=\mu_0H(1+\chi)"

"H=\frac{Ni}{l}"

"H=\frac{220\times1.6}{7}=50.28A/m"

"\chi=\frac{\mu}{\mu_0}-1"

Put value


"\chi=\frac{1.257\times10^{-6}}{4\pi\times10^{-7}}-1=7.96\times10^{-4}"

"B=4\pi\times10^{-7}\times50.28(1+7.96\times10^{-4})"

"B=6.320\times10^{-6}T"